Solve the following problem set.

Problem Set

1. Calculate the mole fraction of the solute in the following solutions.

 a) 100g C2H6O in 100g H2O

 b) 30% HCI solution by mass


2. A solution us prepared by mixing 1g of ethanol ( C2H6O) with 100g water to give a final volume of 101 ml. Calculate the mole fraction for the solute and solvent,and the molarity of ethanol in the solution. 


1
Expert's answer
2021-04-15T03:22:03-0400

Moles C2H6O. = "\frac{100}{46}=2.1739"


Moles of H2O= "\frac{100}{18}=5.5556"


Total moles

"= 5.5556+2.1739=7.7295"

% = "\frac{2.1738}{7.7295}×100=28.12" %